For a standard normal distribution, find the value of z such that: P(0<Z<z) = 0.4901

A. -1.65
B. +2.33
C. -3.33
D. +1.38
Answer #1

We are given here that:

P( 0 < Z < z ) = 0.4901

Therefore P(Z < z) = P(Z < 0) + P(0 < Z < z ) = 0.5 + 0.4901 = 0.9901

From standard normal tables, we have:

P( Z < 2.33) = 0.9901

Therefore +2.33 is the required z value here.